40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4729     Entry time: Tue May 17 01:05:56 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Asymmetry measurement prep : recentering works 

I re-centered beams on several PDs and a camera including :

  AS55, ETMY_QPD, TRY and ETMYT_CCD.

 

The most important one was AS55.

When I was locking each arm I found that the error signal from AS55 was very coupled to the angular motion of the arms.

I checked the beam on the AS55 RFPD and found the beam on the edge of the photo diode. This is possibly because Valera and I had been touching the input beam alignment.

At that time the DC signal from AS55 without aligning PRM and SRM was about 5 mV.

Adjusting the beam position by a steering mirror brought the DC signal up to 20 mV.

Then the lock of each arm became more stable.

ELOG V3.1.3-